### v4

# Magnetic Structures Database from Symmetry-aided High-Throughput Calculations

Abstract
Magnetic structures, which play a central role in determining their physical properties, are known for only very limited compounds. Traditional theoretical approaches to predicting magnetic structures predominantly rely on first-principles calculations. A key challenge of these methods is their requirement for initial magnetic configurations as inputs, which theoretically possess infinite possibilities. In this work, we introduce a strategy based on irreducible representation basis vectors that effectively narrows down the vast space of potential magnetic configurations to a finite set, typically comprising around 20 candidates per material. Despite this significant reduction, the compact input sets generated by our method already encompass the experimental magnetic structures for 253 out of 302 benchmark materials (83.8%) from the MAGNDATA database. These materials have propagation vectors $\mathbf{q}=0$ and unit cells containing up to 40 atoms, all within the Landau framework. Subsequent first-principles calculations correctly identify the magnetic structure in 198 of these cases. We further apply our highly efficient method to 8,422 stoichiometric transition-metal compounds with fewer than 30 atoms per unit cell in the Inorganic Crystal Structure Database, and establish a magnetic structure database containing 2,906 magnetic materials. To demonstrate its utility, we use this database for the systematic exploration of magnetic topological phases and altermagnets, identifying 1,070 and 392 candidate materials, respectively.
I Introduction
Magnetic materials are pivotal to modern technology, enabling applications such as information storage, magnetic topology, spintronics and magnonics through their controllable magnetic response Dagotto et al. (2001); Coey (2020); Bernevig et al. (2022); Eerenstein et al. (2006); Žutić et al. (2004); Han et al. (2020); Chumak et al. (2015); McClarty (2022). They also occupy a central position in condensed matter physics, providing a fertile platform for exploring fundamental collective phenomena that arise from electronic correlations, symmetry breaking, quantum fluctuations, etc. Broholm et al. (2020); Vojta (2018); Löhneysen et al. (2007); Šmejkal et al. (2022a, b); Ma et al. (2021). The essential characteristic of a magnetic material is its magnetic structure, the spatial arrangement of magnetic moments, which determines the magnetic space group and strongly influences its all physical properties. Establishing the magnetic structure is therefore a prerequisite for studying magnetic phenomena.
Despite its importance, the determination of magnetic structures remains a major challenge. Experimentally, magnetic structures can be probed by techniques such as neutron diffraction, yet such methods are often facing technical constrains, including sample quality, crystal size, material-specific factors, and dimensionality Naish et al. (2012); Yankova et al. (2012). As a result, experimentally confirmed magnetic structures are available for only a small fraction of known magnetic compounds Gallego et al. (2016), underscoring the critical role of theoretical prediction.
While symmetry-based frameworks provide a systematic language for describing magnetic configurations Perez-Mato et al. (2024), theoretical prediction of magnetic structures is most commonly carried out using first-principles total-energy calculations. An essential limitation of such approaches is their dependence on an initial set of magnetic configurations as input, which in principle spans an unbounded space. If the experimental magnetic structure is absent from this initial set, first-principles calculations are fundamentally unable to reproduce it. Consequently, the central challenge lies in the absence of an efficient theoretical strategy to generate a finite set of candidate magnetic configurations that can reliably contain the experimental magnetic structure.
Theoretical progress has been ongoing through a variety of strategies Horton et al. (2019); Zheng and Zhang (2021); Tellez-Mora et al. (2024); Sødequist and Olsen (2024); Huebsch et al. (2021); Nomoto et al. (2024), including constrained searches focusing on collinear magnetism Horton et al. (2019), stochastic optimization schemes such as genetic algorithms that explore magnetic configurations without explicit symmetry assumptions Zheng and Zhang (2021), the linear spin wave theory based iterative optimization approach Tellez-Mora et al. (2024), and spin-spiral techniques that efficiently capture long-wavelength noncollinear magnetic orderings within simple magnetic unit cells and have successfully determined 194 magnetic structures Sødequist and Olsen (2024). In addition, Huebsch et al. Huebsch et al. (2021); Nomoto et al. (2024) employed a cluster multipole framework and applied it to the AtomWork-Adv database Xu et al. (2011), which contains 228 magnetic compounds. However, magnetic materials exhibit a remarkable diversity of magnetic orderings, including both collinear and noncollinear states. Moreover, the Inorganic Crystal Structure Database (ICSD) Zagorac et al. (2019) contains on the order of $10^{5}$ materials hosting magnetic ions. These call for a scalable and highly efficient framework capable of treating the full diversity of magnetic orderings, including both collinear and noncollinear states, on an equal footing, irrespective of unit cell complexity.
In this work, we develop a magnetic-configuration strategy based on irreducible-representation (irrep) basis vectors. By construction, the resulting initial configuration space is extremely compact, typically comprising only about twenty candidates per material. Despite this drastic reduction, the generated candidate sets already contain the experimental magnetic structures for the majority of benchmark cases. Namely, for 253 out of 302 materials (83.8%) in the MAGNDATA database with propagation vector $\mathbf{q}=0$ and fewer than 40 atoms per unit cell, the experimental magnetic structures are explicitly included. Subsequent first-principles total-energy calculations then correctly identify the magnetic structure in 198 of these cases, corresponding to accuracy of 78.2%. Applying this approach to compounds with propagation vector $\mathbf{q}=0$ and fewer than 30 atoms per unit cell in ICSD Zagorac et al. (2019), we construct a database comprising over 2,900 magnetic materials. As the magnetic structure fundamentally governs all the physical properties of a material, our magnetic structure database provides a broader materials basis for systematic investigations of physical properties, such as spintronics, magnonics, magnetic topology, multiferroics, etc. Dagotto et al. (2001); Coey (2020); Bernevig et al. (2022); Eerenstein et al. (2006); Žutić et al. (2004); Han et al. (2020); Chumak et al. (2015); McClarty (2022); Broholm et al. (2020); Vojta (2018); Löhneysen et al. (2007); Šmejkal et al. (2022a, b); Ma et al. (2021). As an illustration of its capabilities, we present its application to the systematic identification of magnetic topological phases and altermagnets.
II Method
It is well known that most magnetic phase transitions are continuous and can be described within Landau’s theoryHahn et al. (1983). At the critical temperature, the order parameter emerges continuously from zero, and the resulting magnetic order can be characterized by the irreducible representation of the parent crystallographic space group. This motivates us to construct a set of symmetry-guided magnetic structure candidates corresponding its irreducible representation. Our strategy is as follows.
For a continuous phase transition, the magnetic space groups (MSGs) of the ordered phase must be a subgroup of the parent crystallographic space group. However, the subgroup landscape is typically vast. To systematically reduce the search space of candidate MSGs compared with a brute-force subgroup search, we take the magnetic moments $\boldsymbol{\mu}$ on the magnetic Wyckoff positions as a basis. For a crystal containing $N$ magnetic ions in the unit cell, magnetic configurations with propagation vector $\mathbf{q}=0$ are described by the $3N$ components of the local magnetic moments, $\boldsymbol{\mu}=(m_{1x},m_{1y},m_{1z},\dots,m_{Nx},m_{Ny},m_{Nz})$, where $m_{na}$ denotes the magnetic-moment component of atom $n$ along the crystallographic axis $a$. The $3N$ components of $\boldsymbol{\mu}$ provide a basis for representation $\Gamma_{\mathrm{mag}}$, which is generally reducible and can be decomposed into irreducible representations as
$\Gamma_{\mathrm{mag}}=\sum_{\oplus i}n_{i}\,\Gamma_{i}(j).$ (1)
Here, $\Gamma_{i}(j)$ denotes an irreducible representation of dimension $j$ from the parent crystallographic space group, and $n_{i}$ is its multiplicity. The completeness relation $\sum_{i}n_{i}\cdot j=3N$ ensures that all magnetic degrees of freedom are fully accounted for in this decomposition.
Within Landau theory, a continuous magnetic transition is driven by the instability of a single irreducible representation in Eq. (1). Consequently, only the MSGs compatible with the unstable irrep, namely those that actually appear in Eq. (1), need to be considered. This restriction substantially reduces the search space of candidate MSGs. Even after the unstable irrep and its compatible MSGs are fixed, the magnetic structure is, in general, still not uniquely determined. To construct explicit magnetic configurations, one needs to further specify the symmetry-adapted basis vectors of the irreducible representation. However, even within this basis-vector description, residual degrees of freedom may remain. This non-uniqueness originates from the fact that for multidimensional irreps ($j>1$) and/or repeated irreps ($n_{i}>1$) in Eq. (1), the basis vectors are not uniquely fixed and can be linearly combined in symmetry-allowed ways. In the following, we describe how the associated degrees of freedom can be systematically constrained.
For multi-dimensional irreducible representations $\Gamma_{i}(j)$ with $j>1$ in Eq. (1), the set of the order parameter within the irrep space is not uniquely determined, corresponding to different MSGs. Although, in principle, the physically realized order parameter could be obtained by minimizing higher-order terms of the Landau free energy, this route is not adopted in the present work. Instead, we adopt a symmetry-based criterion and retain all the maximal-symmetry magnetic subgroups among those compatible with the irrep in Eq. (1). Physically, ordering patterns belonging to these MSGs correspond to minimal symmetry breaking at the magnetic transition Perez-Mato et al. (2015); Zhou et al. (2023).
A second source of non-uniqueness arises when a given irreducible representation appears multiple times in the representation, i.e., $n_{i}>1$ in Eq. (1). In this case, the magnetic structure depends on the particular choice of basis, since multiple symmetry-equivalent sets of basis vectors exist and can be linearly combined, giving rise to an infinite number of possible initial input magnetic configurations. To construct a finite set of magnetic configurations, we keep only irreducible-representation basis vectors, whose resulting magnetic moments align with all possible high-symmetry crystallographic directions Hahn et al. (1983) (see Supplementary Materials SM ).
For each candidate material, a finite set of usually $\sim$20 symmetry-allowed magnetic structures associated with the relevant irreducible representations was generated using the procedure described above. First-principles total-energy calculations were then performed among these well-selected magnetic structures using the Vienna Ab initio Simulation Package (VASP) Kresse and Joubert (1999); Kresse and Furthmüller (1996), employing a spin-polarized GGA + SOC + $U$ approach. The Coulomb interaction parameters $U$ and $J$ were taken as the averaged values recommended for transition-metal elements based on the linear-response scheme Moore et al. (2024). The ground-state magnetic structure was identified as the configuration with the lowest total energy.
We benchmark our theory framework against the MAGNDATA database Gallego et al. (2016), focusing on stoichiometric materials with fewer than 40 atoms per unit cell and magnetic propagation vector $\mathbf{q}=0$. Only compounds containing transition-metal ions are considered; lanthanoid and actinoid systems are excluded due to the difficulty of reliably treating within standard density functional theory Kotliar et al. (2006). These criteria yield a total of 328 materials, among which 26 are excluded because their experimental magnetic structures are incompatible with continuous magnetic phase transitions (see Supplementary Materials SM ).
For the remaining 302 materials, our irrep basis-vector strategy systematically reduces the vast space of initial magnetic configurations to a finite set of typically $\sim$20 candidates per material. Remarkably, the experimental magnetic structures of 253 materials (83.8%) are already contained within these candidate sets. Using the generated initial input magnetic configurations ($\sim 5000$ for 253 materials in total), subsequent first-principles total-energy calculations correctly identify the magnetic structures for 198 cases, corresponding to an accuracy of 78.2%.
The noncollinear magnet Mn3GaN Nan et al. (2020) is presented in the Supplementary Materials SM to illustrate our complete workflow. The crystallographic space group of Mn3GaN yields 98 magnetic subgroups and 10 inequivalent irreducible representations. By incorporating the Wyckoff positions of Mn ions, these 10 irreducible representations are reduced to only two relevant ones, $\Gamma_{4}^{+}$ and $\Gamma_{5}^{+}$, together with 9 symmetry-compatible magnetic space groups. Applying the maximal-symmetry criterion further narrows this set to 5 maximal magnetic subgroups. Finally, for the case of repeated irreducible representations ($n_{i}>1$ in Eq. (1)), we retain only basis vectors that generate magnetic moments aligned along high-symmetry crystallographic directions. As a result, the originally vast configuration space is reduced to just 11 candidate magnetic structures, within which the experimental magnetic structure of Mn3GaN is included and subsequently identified by first-principles calculations.
Figure 1: Magnetic and topological classification of our magnetic structure database. (a) Magnetic structure classification. Among the 2,906 identified magnetic materials, 49.86% exhibit ferromagnetic (FM) ordering, 11.63% are ferrimagnetic (FiM), and the remaining 38.51% are antiferromagnetic (AFM). The FiM materials are further classified into 7.67% collinear (Col) and 3.96% noncollinear (NoC) states according to the relative orientation of magnetic moments. Within the AFM materials, 21.54% exhibit collinear (Col) order, 3.48% exhibit noncollinear (NoC) order, and 13.49% are classified as altermagnets (Alter). (b) Topological classification of magnetic materials. In total, 38% of the magnetic materials in the database are identified as magnetic topological materials, including 4% magnetic topological insulators (TIs) and 34% magnetic topological semimetals (TSMs).
Motivated by the excellent benchmark performance, we applied our framework to the Inorganic Crystal Structure Database (ICSD) Zagorac et al. (2019). Focusing on stoichiometric transition-metal compounds with fewer than 30 atoms per unit cell and excluding lanthanoids and actinoids, we restricted our analysis to magnetic structures with propagation vector $\mathbf{q}=0$. Within this well-defined scope, our irrep basis-vector strategy generates a compact set of trial magnetic configurations, typically <20<20 per material, enabling systematic high-throughput first-principles calculations. Out of 8,422 compounds examined, with a total of approximately 170,000 generated initial magnetic configurations, 5,516 materials are identified as non-magnetic. The remaining 2,906 magnetic materials, together with their magnetic structures, magCIF files, and associated metadata, have been compiled into a searchable database, available at https://magdatabase.nju.edu.cn.
This database contains a broad diversity of magnetic configurations. Among the identified magnetic structures, 2690 are collinear and 216 are noncollinear. From the viewpoint of magnetic classification, 1,449 materials exhibit ferromagnetic (FM) ordering, and 338 materials are classified as ferrimagnetic (FiM). 1,119 compounds exhibit antiferromagnetic (AFM) order with vanishing net magnetization, including 392 altermagnets. The classification statistics of the magnetic structures are shown in Fig. 1. A complete list of the magnetic materials classification statistics is presented in the Supplementary Materials SM .
Beyond serving as a catalog of magnetic structures, we further apply the database to perform topological classification and altermagnets, as will be demonstrated below.
III Magnetic Topological Materials
Figure 2: AFM axion insulator FeSe. (a) magnetic structure (b) band structure. The green and yellow balls are Se and Fe ions, respectively. The magnetic moments of Fe are along the $b$-axis ([010]), as indicated by red arrow.
For decades, the topology of electronic bands in condensed-matter materials has been a research hotspot Hasan and Kane (2010), leading to a systematic classification of several tens of thousands non-magnetic topological materials based on the symmetry analysis Zhang et al. (2019); Vergniory et al. (2019); Tang et al. (2019). Owing to the limited number of known magnetic structures Gallego et al. (2016), theoretical predictions of magnetic topological materials remain relatively scarce, with only a few hundred reported to date Xu et al. (2020); Bernevig et al. (2022). Based on our magnetic structure database, we identify 1070 compounds as magnetic topological materials using magnetic topological quantum chemistry Kruthoff et al. (2017); Elcoro et al. (2021). They can be broadly classified into 125 magnetic topological insulators and 945 magnetic topological semimetals, as summarized in Fig. 1b. A complete list of materials and symmetry indicators is provided in the Supplementary Materials SM . To illustrate the physical richness enabled by the symmetry-resolved magnetic structures, we select several representative materials shown below. We also tested the effects of varying the Hubbard $U$ within reasonable ranges, and found that both their magnetic structures and topological properties remain unchanged.
III.1 Magnetic topological insulators
AFM axion insulator FeSe: The FeSe compound (icsd_166445) studied here crystallizes in the orthorhombic space group $Pnma$ (No. 62) Margadonna et al. (2009), which is different from the widely studied superconducting $\alpha$-FeSe of the tetragonal structure $P4/nmm$ (No. 129) Zhang et al. (2018); Wang et al. (2018). Our calculations predict an antiferromagnetic ground state in FeSe. It belongs to the magnetic space group $Pnma$ (62.441), with magnetic moments oriented along the [010] direction, as shown in Fig. 2a. Due to the presence of inversion symmetry, symmetry indicators ($\eta_{4I}$, $z_{2I,1}$, $z_{2I,2}$, $z_{2I,3}$) can be defined using the parities at high-symmetry points Elcoro et al. (2021). $\eta_{4I}$ is defined as the total number of odd-parity states at the eight time-reversal-invariant momenta (TRIMs) modulo 4, while $z_{2I,i}$ is defined as the total number of odd-parity states at the four TRIMs on the $k_{i}=\pi$ plane modulo 2. Our calculated symmetry indicators ($\eta_{4I}$, $z_{2I,1}$, $z_{2I,2}$, $z_{2I,3}$) are (2000). Furthermore, we find that all plane Chern numbers on the TRIM planes vanish for FeSe, and the system exhibits a direct band gap larger than 20 meV throughout the entire Brillouin zone, as shown in Fig. 2b. The nontrivial symmetry indicators, together with a direct band gap throughout the entire Brillouin zone, indicate that FeSe is an axion insulator, in which a topological magnetoelectric effect is expected to occur Essin et al. (2009); Wan et al. (2012).
III.2 Magnetic topological semimetals
Figure 3: Topological magnetic semimetals. (a)-(c) Magnetic structure, band structure, and distribution of Weyl points of the non-collinear AFM Weyl semimetal V3Ga2. The magnetic moments of V lie in the xyxy plane, as indicated by red arrow. (d)-(f) Magnetic structure, band structure, and distribution of Dirac points of the AFM Dirac semimetal MnTe. (g)-(i) Magnetic structure, band structure, and distribution of nodal rings and Weyl points of the FM nodal-line semimetal Co3SnC. The red and blue dots in (c) and (k) denote Weyl points with topological charges +1 and 1-1, respectively, while the green lines represent the nodal rings.
[…]
AFM Dirac semimetal MnTe: The paramagnetic phase of MnTe (icsd_076241) has the space group Pnma (No. 62), which is the high-pressure structure of altermagnet $\alpha$-MnTe Krempaský et al. (2024); Lee et al. (2024). Under high pressure, $\alpha$-MnTe undergoes a structural phase transition from the NiAs-type structure to the MnP-type structure Mimasaka et al. (1987). We find that the magnetic ground state of MnTe (icsd_076241) is a $PT$ antiferromagnet, belonging to the magnetic space group $Pn^{\prime}m^{\prime}a^{\prime}$ (62.449), as shown in Fig. 3d. Throughout the entire Brillouin zone, two Dirac points are located on the high-symmetry line $k_{y}=0$, $k_{z}=\pi$, while four Dirac points reside on the high-symmetry line $k_{y}=\pi$, $k_{z}=0$, as shown in Fig. 3e and f.
FM nodal-line semimetal Co3SnC: The Co3SnC (icsd_108129) adopts an anti-perovskite structure with the $Pm\bar{3}m$ (No. 221) space group Lin et al. (2014). Its magnetic ground state is a ferromagnet with magnetization oriented along the $[001]$ direction. The magnetic structure and the calculated electronic band structure of Co3SnC are shown in Fig. 3g and 3h. As shown in the Fig. 3i, two nodal rings appear on the high-symmetry planes $k_{z}=0$ and $k_{z}=\pi$. The two crossing bands carry +i+i and i-i eigenvalues of the mirror symmetry $m_{z}$, respectively, and are thus protected by $m_{z}$ symmetry. In addition, 16 Weyl points are present in the Brillouin zone, with energies ranging from 0.1-0.1 to $0.15$ eV.
Figure 4: (a) Crystal structure and (b) band structure of CsV2S2O.
IV Altermagnets
Altermagnets represent a recently identified magnetic class distinct from conventional antiferromagnets, garnering significant attention due to their unique electronic band structures, chiral magnon splitting, etc. Šmejkal et al. (2022a, b); Xu et al. (2025).
In our database, we have identified 392 altermagnetic materials, many of which exhibit large electronic spin splitting and chiral magnon splitting, with a complete list provided in the Supplementary Materials SM . Here we highlight a representative example, CsV2S2O.
CsV2S2O (icsd_430299) crystallizes in space group $P4/mmm$ (No. 123), and its magnetic moments are aligned along [001] direction, see Fig. 4(a). It exhibits strong altermagnetic spin splitting along high-symmetry paths [Fig. 4(b)], reaching a maximum of approximately 1.7 eV near the X point, indicative of a pronounced electronic altermagnetic effect. The magnon spectrum shows a chiral splitting of up to 94.4 meV at the X point in the Brillouin zone, originating from two symmetry-inequivalent next-nearest-neighbor exchange interactions SM , in contrast to the typical origin which arises from disparities in long-range exchange interactions (e.g., $J_{7a}$ versus $J_{7b}$ in MnF2 Morano et al. (2025)). This value significantly exceeds the experimental chiral magnon splittings in MnF2 Morano et al. (2025) and MnTe Liu et al. (2024).
V Outlook
We have developed a highly efficient methodology that addresses a bottleneck in magnetic materials research by combinating symmetry analysis with first-principles calculations. For continuous magnetic phase transitions, the symmetry of the ordered phase follows a subgroup relation with respect to the parent space group. Using the Wyckoff positions of magnetic atoms, we construct the magnetic representation and decompose it into the irreducible representations of the parent crystallographic space group, thereby reducing the space of candidate MSGs to those compatible with the corresponding irreducible representations in Eq. (1). As the irreps and the corresponding MSGs do not uniquely fix the magnetic structure, we employ the symmetry-adapted basis vectors of the irreducible representation, which provide a more proper strategy for constructing explicit magnetic configurations, yet residual degrees of freedom remain. We further resolve these ambiguities by applying the maximal-symmetry criterion, and by keeping all the basis vectors of irreducible representations compatible with all possible highest-symmetry crystallographic directions. In this way, our strategy yields a group theory based framework that drastically reduces the initial magnetic configuration space, usually $\sim 20$ candidates, and thereby enables efficient high-throughput first-principles calculations.
To assess the scope of our irreducible representation basis vector based strategy, we analyzed 302 materials in the MAGNDATA database. For all materials whose magnetic structures arise from multidimensional irreducible representations ($j>1$ in Eq. (1)), the experimental magnetic structures are fully captured by the maximal-symmetry criterion. By construction, our strategy reduces the otherwise large space of magnetic configurations to compact candidate sets, typically consisting of only tens of predefined configurations per material. Remarkably, these minimal input sets already contain the experimental magnetic structures for 253 compounds (83.8%; see Supplementary Materials SM ). The remaining cases fall outside the scope of our framework, as their magnetic moments deviate from high-symmetry crystallographic directions. Starting from this strongly reduced configuration space, first-principles calculations correctly identify the magnetic structure in 198 of the 253 materials, corresponding to an accuracy of 78.2%, which is expected to further improve with advances in electronic-structure methods.
The efficiency and scalability of our framework allow the construction of a comprehensive magnetic-structure database. In the present work, we focus on compounds with propagation vector $\mathbf{q}=0$ and fewer than 30 atoms per primitive cell as a starting point. Extensions to larger unit cells and to magnetic structures with $\mathbf{q}\neq 0$ are straightforward within our framework and are currently underway. Moreover, applying our strategy to two-dimensional systems is particularly timely and significant, given the rapidly growing interest in low-dimensional magnetism. Beyond the topological classification already established in this database, we are actively exploring additional magnetic-related physical properties, including multiferroic responses, anomalous Hall conductivity, transport coefficients, etc.
VI Acknowledgements
This work was financially supported by the National Natural Science Foundation of China (Grants No. 12188101, No. 12574066, No. 12474233, and No. 12334007), the National Key R&D Program of China (Grant No. 2025YFA1411301, No. 2022YFA1403601), Innovation Program for Quantum Science and Technology (Grant No.2021ZD0301902), Quantum Science and Technology-National Science and Technology Major Project (No. 2024ZD0300101), Natural Science Foundation of Jiangsu Province (No. BK20233001, BK20243011), Fundamental and Interdisciplinary Disciplines Breakthrough Plan of the Ministry of Education of China (JYB2025XDXM411), the Fundamental Research Funds for the Central Universities (No. KG202501). This work has been supported by the New Cornerstone Science Foundation.